OpenStudy (anonymous):

The function f(t) = 4 cos(pi over 3t) + 15 represents the tide in Bright Sea. It has a maximum of 19 feet when time (t) is 0 and a minimum of 11 feet. The sea repeats this cycle every 6 hours. After five hours, how high is the tide? 13.5 feet 16 feet 18.5 feet 17 feet

OpenStudy (phi):

you need to find the value of cos(5pi/3) in degrees that is 5*180/3 = 5*60 = 300º 300º is in the 4th quadrant. it will be the same value as cos(60º)

OpenStudy (phi):

\[ f(5) = 4\cos\left( \frac{\pi}{3} \cdot 5 \right)+15 \]

OpenStudy (anonymous):

sorry my open study is messing up right now

OpenStudy (anonymous):

omg my calculator just died.. can u plug that in and see what u get?

OpenStudy (phi):

first, people memorize the sin, cos and tan of certain angles: 0,30,45,60 and 90 so you should know 60º second, type it into the google search window (but google uses radians)

OpenStudy (anonymous):

i got 16 feet

OpenStudy (phi):

what did you get for the cos part?

OpenStudy (anonymous):

16

OpenStudy (phi):

cos is between -1 and 1 what is cos(5 pi/3) ?

OpenStudy (anonymous):

its not working

OpenStudy (phi):

type into google cos(5 pi/3)=

OpenStudy (anonymous):

.5

OpenStudy (phi):

now do f(5)= 4*0.5 + 15

OpenStudy (anonymous):

17

OpenStudy (phi):

yes

OpenStudy (anonymous):

thank u